Consider Higgins Production which has the following information about its capital structures: Debt - 4,500, 5% coupon bonds outstanding, $1,000 par value, 7 years to maturity, selling for 80 % of par, the bonds make semiannual payments • Common Stock - 100,000 shares outstanding, selling for $35 per share; the beta is 1.20 • Preferred Stock - 19,000 shares of 6 % preferred stock outstanding, currently selling for $150 per share • Market Information - 6 %t market risk premium and 4 % risk-free rate. Required: Calculate to the following if the company has a tax rate of 36 percent. i. Total Market Value for the Firm ii. After-tax cost of Debt iii. Cost of Equity iv. Cost of Preferred Stock v. Weighted Average Cost of Capital
Answer #1
MV of equity=Price of equity*number of shares outstanding
MV of equity=35*100000
=3500000
MV of Bond=Par value*bonds outstanding*%age of par
MV of Bond=1000*4500*0.8
=3600000
MV of Preferred equity=Price*number of shares outstanding
MV of Preferred equity=150*19000
=2850000
i) MV of firm = MV of Equity + MV of Bond+ MV of Preferred equity
=3500000+3600000+2850000
=9950000
Weight of equity = MV of Equity/MV of firm
Weight of equity = 3500000/9950000
W(E)=0.3518
Weight of debt = MV of Bond/MV of firm
Weight of debt = 3600000/9950000
W(D)=0.3618
Weight of preferred equity = MV of preferred equity/MV of firm
Weight of preferred equity = 2850000/9950000
W(PE)=0.2864
III Cost of equity
As per CAPM
Cost of equity = risk-free rate + beta * (Market risk premium)
Cost of equity% = 4 + 1.2 * (6)
Cost of equity% = 11.2
Cost of debt
                  K = Nx2
Bond Price =∑ [(Semi Annual Coupon)/(1 + YTM/2)^k]     +   Par value/(1 + YTM/2)^Nx2
                   k=1
                  K =7x2
800 =∑ [(5*1000/200)/(1 + YTM/200)^k]     +   1000/(1 + YTM/200)^7x2
                   k=1
YTM = 8.900128943
ii. After tax cost of debt = cost of debt*(1-tax rate)
After tax cost of debt = 8.900128943*(1-0.36)
= 5.70
Iv. cost of preferred equity
cost of preferred equity = Preferred dividend/price*100
cost of preferred equity = 6/(150)*100
=4
v. WACC=after tax cost of debt*W(D)+cost of equity*W(E)+Cost of preferred equity*W(PE)
WACC=5.7*0.3618+11.2*0.3518+4*0.2864
WACC =7.15%